N2(g) + 3 H2(g) ⇌ 2 NH3(g) KP = 6.78 x 105 at 298 K (determined using atm)

A 7.5 x 101 L container being held at 298 K is charged with the three gases present in the above equation. Once finished, the initial partial pressure of N2 was 0.59 atm, the initial partial pressure of H2 was 0.45 atm, and the initial partial pressure of NH3 was 0.11 atm. The gas mixture was then allowed to reach equilibrium. Use this information to determine the partial pressures for all three gasses after equilibrium is reached.
Answer #1

Attached below is the handwritten solution to the above-mentioned problem.

N2 + 3 H2 = 2NH, ; Kp = 6.78x10 Step 1: Determine the reaction quotient) - Q= PNike (0.11² - 0.255< ke PNX PH (0.59)*(0:45) e

. We know that value of x will never be greater than either of the reactants. is. x<0.59 and 3 x < 0.45= 2<.0.15 => 1C 20.15

- - On good further enough guessing we see that a = 0.146 is approximation for a. a 1. Equilibrium come are: PN = 0.59-0.146

On Solving we get the partial pressure of the gases at equilibrium as: (Approximately)

PN2 = 0.444atm

PH2 = 0.406atm

PNH3 = 0.139atm